# A Linear Kernel for Planar Vector Domination

### v1

@@ Abstract
Given a graph $G$, an integer $k\geq 0$, and a non-negative integral function $f:V(G)\rightarrow\mathcal{N}$, the Vector Domination problem asks whether a set $S$ of vertices, of cardinality $k$ or less, exists in $G$ so that every vertex $v\in V(G)-S$ has at least $f(v)$ neighbors in $S$. The problem generalizes several domination problems and it has also been shown to generalize Bounded-Degree Vertex Deletion. In this paper, the parameterized version of Vector Domination is studied when the input graph is planar. A linear problem kernel is presented.
@@ 1 Introduction
The Vector Dominating Set problem assumes a graph $G$ is given along with an integer $k\geq 0$ and a demand function $f$ from the set $V$ of vertices of $G$ to the set $\{0,\ldots|V|-1\}$, and asks whether a set $D$ of at most $k$ vertices of $G$ exists such that each vertex $v$ has at least $f(v)$ neighbors in $D$. The problem can be viewed as a generalization of many domination problems that are extensively studied in the literature, including the classic Dominating Set, the $k$-dominating set problem [18, 4, 5, 6, 9, 15] and the bounded degree deletion problem [12, 3, 10] along with its dual problem the $s$-Plex Detection problem, which finds application in social network analysis [20, 2].
@@
From a computational complexity standpoint, Vector Dominating Set is $NP$-hard being a generalization of the dominating set problem. Thus it is natural to ask under what conditions the problem becomes tractable. It is well known by now that the study of a problem’s parameterized complexity [7, 8] provides an alternative view of its complexity and can shed light on classes/types of input where the problem can be tractable. Under the parameterized complexity lens, a problem is fixed parameter tractable (FPT) if there exists an algorithm that solves it in $O(f(k)n^{c})$ time where $n$ is the size of the input, $k$ is an input parameter, $f$ is a function of $k$ and $c$ is a constant. Alternatively, a problem is FPT if it admits a kernel i.e given an instance $(G,k)$ of the problem, an equivalent instance $(G^{\prime},k^{\prime})$ can be constructed in polynomial time where $|G^{\prime}|\leq g(k)$ and $k^{\prime}\leq k$. $G^{\prime}$ is referred to as a problem kernel. Of particular interest is the case where $g$ is a (low degree) polynomial function of $k$. The dominating set problem is $W[2]$-hard on general graphs [11], thus there is little or no hope in achieving fixed parameter tractability for it or for the more general vector dominating set problem on general graphs.
@@
Although planar vector domination is known to be fixed-parameter tractable [17], a natural question to ask is whether the problem admits a polynomial kernel on classes of graphs where it is known to be fixed-parameter tractable. On planar graphs, Alber et al. (2004) provided an explicit linear kernel for the dominating set problem. This was based on structural properties of planar graphs, especially degree structure, and it was followed by a series of papers on linear kernels of different domination problems [1, 14, 13, 19]. No linear kernel is known yet for Vector Dominating Set in planar graphs. Addressing this problem is the main objective of this paper. A linear kernel is presented based on several structural properties of planar graphs and the approach of Alber et al.. The bound achieved is $101k$, which is comparable to the upper bounds for related problems. Moreover, the obtained bound leads to a linear kernel for the Planar Bounded-Degree Vertex Deletion problem.
@@ 3 A Linear Kernel
From this point one, we denote by $A$ the set of 0-vertices such that $N(A)\subset\{u\in V(G):|N(u)\cap S|\geq 2\}$. Set $L$ to be the set of all leeches for real paths.
@@
Suppose V(G) contains a solution set $S$, then after applying the reduction rules above any vertex $v\in G-(S\cup A\cup L)$ belongs to some real path.
@@
Now suppose that $d(v)=0$. If $|N(v)\cap S|\geq 2$, $v$ lies on a type 1 real path, and as above if $N(v)\cap S=\{s_{1}\}$ then there exists a real path $s_{1}vus_{2}$. So we may suppose that $N(v)\cap S=\emptyset$. Since $v\notin A$, there exists $u\in N(v)$ where $N(u)\cap S=\{s_{1}\}$, As above using reduction rule 4 we may verify that there exists a real path $s_{1}uvu^{\prime}s_{2}$. ∎
@@
In the following rules let $a_{1},a_{2}\in V(G)$,$C(a_{1},a_{2})$ be a candidate region between $a_{1}$ and $a_{2}$. Let $C,O,B,I,I^{\prime}$ and $O^{\prime}$ denote $C(a_{1},a_{2})$, $O(a_{1},a_{2})$, $B(a_{1},a_{2}),$ $I(a_{1},a_{2}),$ $I^{\prime}(a_{1},a_{2})$ and $O^{\prime}(a_{1},a_{2})$ respectively. Reduction rules $6,7$ and $8$ apply only on planar graphs. They are not applied exhaustively.
@@
Let $G$ be a planar graph, then computing a vector dominating set for $G$ of size at most $k$ can be reduced to computing a vector dominating set of size at most $k$ to a planar graph $G^{\prime}$ where $|G^{\prime}|\leq 101k$.
@@
Claim1: Each vertex of $A_{1}\cup L$ has three or more neighbors in $K$. To see this, take $c\in A_{1}$ then $N(c)\cap L=\emptyset$ (a leech cannot be adjacent to a 0-vertex) so $\forall v\in N(c)$, $v\in\mathscr{R}$ thus $v$ is on some outer path by planarity and since $|N(c)|\geq 3$ Claim1 follows. Now take $l\in L^{\prime}$. $l$ is a leech for some type 3 path, so $\exists s,v\in N(l)$ with $s\in S$ and $d(v)=1$. $v$ must be on some outer path since $v\in V(\mathscr{R})$ ($v\notin A\cup L)$ and $v$ cannot be in the interior of a region due to planarity. Now, by reduction rule 11, $l$ must have a neighbor $v^{\prime}$ different from $s$ and $v$. We must have $d(v^{\prime})=1$ since $l$ is a leech and $v^{\prime}\notin L$ since vertices of $L$ are colored blue by reduction rule 9, so no two leeches can be adjacent. Thus similar to $v$, $v^{\prime}$ is on some outer path. This proves our claim.
@@
Thus we have $|V(G^{\prime})|\leq 15\cdot 3k+(x+3y+2z)+(3x+3z)+26k=71k+4x+3y+5z$ but this expression is maximized if $z=6k$, which gives $|V(G^{\prime})|\leq 101k$. ∎

### v2

@@ Abstract
Given a graph $G$, an integer $k\geq 0$, and a non-negative integral function $f:V(G)\rightarrow\mathcal{N}$, the Vector Domination problem asks whether a set $S$ of vertices, of cardinality $k$ or less, exists in $G$ so that every vertex $v\in V(G)\setminus S$ has at least $f(v)$ neighbors in $S$. The problem generalizes several domination problems and it has also been shown to generalize Bounded-Degree Vertex Deletion (BDVD). In this paper, the parameterized version of Vector Domination is studied when the input graph is planar. A linear problem kernel is presented. A direct consequence is a kernel bound for BDVD that is linear in the parameter $k$ only. Previously known bounds are functions of both the target degree and the input parameter.
@@ 1 Introduction
The Vector Dominating Set problem assumes a graph $G$ is given along with an integer $k\geq 0$ and a demand function $f$ from the set $V$ of vertices of $G$ to the set $\{0,\ldots|V|-1\}$, and asks whether a set $D$ of at most $k$ vertices of $G$ exists such that each vertex $v\in V(G)\setminus D$ has at least $f(v)$ neighbors in $D$. The problem can be viewed as a generalization of many domination problems that are extensively studied in the literature, including the classic Dominating Set problem, the $r$-Dominating Set problem [25, 5, 6, 14, 21] and the $\alpha$-Dominating set problem [13, 8, 9, 17] which is also known to be a generalization of Positive Influence Dominating Set (PIDS). The latter problem is known for applications in social networks [10, 11, 28, 1]. PIDS is also called a monopoly in the literature [24]. In addition to domination problems, Vector Domination is also a generalization of the Bounded Degree Vertex Deletion problem (BDVD) [18, 4, 15], along with its dual, the $s$-Plex Detection problem. The latter also finds application in social network analysis [27, 3].
@@
From a computational complexity standpoint, Vector Dominating Set is $NP$-hard being a generalization of the dominating set problem. It is well known by now that the study of a problem’s parameterized complexity [7, 12] provides an alternative view of its complexity and can shed light on classes/types of input where the problem can be tractable. Under the parameterized complexity lens, a problem is fixed parameter tractable (FPT) if there exists an algorithm that solves it in $O(f(k)n^{c})$ time where $n$ is the size of the input, $k$ is an input parameter, $f$ is a function of $k$ and $c$ is a constant. Alternatively, a problem is FPT if it admits a kernel i.e given an instance $(G,k)$ of the problem, an equivalent instance $(G^{\prime},k^{\prime})$ can be constructed in polynomial time where $|G^{\prime}|\leq g(k)$ and $k^{\prime}\leq k$. $G^{\prime}$ is referred to as a problem kernel. Of particular interest is the case where $g$ is a (low degree) polynomial function of $k$. The Dominating Set problem is $W[2]$-hard on general graphs [16], thus there is little or no hope in achieving fixed parameter tractability for it or for the more general Vector Dominating set problem in general graphs.
@@
On the other hand, Vector Dominating Set is known to be fixed-parameter tractable when the input is restricted to planar graphs [23]. By exploiting structural properties of planar graphs, especially degree structure, Alber et al. [2] provided a linear kernel for the planar Dominating Set problem. This early work was followed by a series of papers on linear kernels for different domination problems [2, 20, 19, 26]. Surprisingly, no linear kernel is known yet for the (more general) Vector Dominating Set problem in planar graphs. Addressing this problem is the main objective of this paper. A linear kernel is presented based on several structural properties of planar graphs, including the methods used in [2]. The bound achieved is $101k$, and it results in linear kernel bounds for several special problems including the Planar Bounded-Degree Vertex Deletion (PBDVD) for which the obtained kernel is linear in the number of deleted vertices ($k$) only. Previously known kernel bounds (in general) are functions of $k$ and the target degree bound [?].
@@ 3 A Linear Kernel
From this point on, we denote by $A$ the set of 0-vertices such that $N(A)\subset\{u\in V(G):|N(u)\cap S|\geq 2\}$. Set $L$ to be the set of all leeches for real paths.
@@
If V(G) contains a solution set $S$, then after applying the reduction rules above any vertex $v\in G-(S\cup A\cup L)$ belongs to some real path.
@@
Now suppose that $d(v)=0$. If $|N(v)\cap S|\geq 2$, $v$ lies on a type 1 real path, and as above if $N(v)\cap S=\{s_{1}\}$ then there exists a real path $s_{1}vus_{2}$. So we may suppose that $N(v)\cap S=\emptyset$. Since $v\notin A$, there exists $u\in N(v)$ where $N(u)\cap S=\{s_{1}\}$, As above using reduction rule 4 we may verify that there exists a real path $s_{1}uvu^{\prime}s_{2}$.
@@
In the following rules let $a_{1},a_{2}\in V(G)$, $C(a_{1},a_{2})$ be a candidate region between $a_{1}$ and $a_{2}$. Let $C,O,B,I,I^{\prime}$ and $O^{\prime}$ denote $C(a_{1},a_{2})$, $O(a_{1},a_{2})$, $B(a_{1},a_{2}),$ $I(a_{1},a_{2}),$ $I^{\prime}(a_{1},a_{2})$ and $O^{\prime}(a_{1},a_{2})$ respectively. Reduction rules $6,7$ and $8$ apply only on planar graphs.
@@
Let $G$ be a planar graph, then computing a vector dominating set for $G$ of order at most $k$ can be reduced to computing a vector dominating set of order at most $k$ to a planar graph $G^{\prime}$ where $|G^{\prime}|\leq 101k$.
@@
Claim1: Each vertex of $A_{1}\cup L$ has three or more neighbors in $K$. To see this, take $c\in A_{1}$ then $N(c)\cap L=\emptyset$ (a leech cannot be adjacent to a 0-vertex) so $\forall v\in N(c)$, $v\in\mathscr{R}$ thus $v$ is on some outer path by planarity and since $|N(c)|\geq 3$ Claim1 follows.
Now take $l\in L^{\prime}$. $l$ is a leech for some type 3 path, so $\exists s,v\in N(l)$ with $s\in S$ and $d(v)=1$. $v$ must be on some outer path since $v\in V(\mathscr{R})$ ($v\notin A\cup L)$ and $v$ cannot be in the interior of a region due to planarity. Now, by reduction rule 11, $l$ must have a neighbor $v^{\prime}$ different from $s$ and $v$. We must have $d(v^{\prime})=1$ since $l$ is a leech and $v^{\prime}\notin L$ since vertices of $L$ are colored blue by reduction rule 9, so no two leeches can be adjacent. Thus similar to $v$, $v^{\prime}$ is on some outer path. This proves our claim.
@@
Thus we have $|V(G^{\prime})|\leq 15\cdot 3k+(x+3y+2z)+(3x+3z)+26k=71k+4x+3y+5z$ but this expression is maximized if $z=6k$, which gives $|V(G^{\prime})|\leq 101k$.
The above bound applies to several problems that can be seen as special cases. In particular, the same kernel bound is obtained for Bounded Degree Vertex Deletion.
Corollary 1
The Planar Bounded Degree Vertex Deletion problem admits a kernel of order at most $101k$ irrespective of the target degree bound.
The same applies for the $r$-Dominating Set problem, which asks, for a given planar graph $G=(V,E)$ and integers $r,k$, whether there is a set $D$ of at most $k$ vertices such that every element of $V\setminus D$ has at least $r$ neighbors in $D$. Obviously, this is a special case of Vector Domination where the demand vector $d$ satisfies $d(v)=r,\forall v\in V$.
Corollary 2
The Planar $r$-Dominating Set problem admits a kernel of order at most $101k$.
Another well-known special case of Vector Domination is the $\alpha$-Dominating Set problem. Given a graph $G=(V,E)$ along with an integer $k$ and $\alpha\in(0,1$], the $\alpha$-domination problem asks for a set $D\subset V$ of cardinality at most $k$ such that every vertex $w\in V\setminus D$ has a at least $\alpha\times degree(w)$ neighbors in $D$.
Corollary 3
The $\alpha$-Dominating Set problem admits a kernel of order at most $101k$.
The same kernel bound is also obtained for the Positive Influence Dominating Set problem (PIDS), which (typically) corresponds to the case where $\alpha=0.5$.